Solve this system of equations using the elimination method:
3a-3b=-15
-3a-3b=-3
please explain steps
thank you

Answer by jim_thompson5910(28476)

Simply add the equations to get

3a-3b=-15
-3a-3b=-3
---------
0a-6b=-18

So

-6b = -18

b = -18/(-6)

b = 3

Now that we know that b = 3, we can use this to find 'a'.

3a-3b=-15

3a-3(3)=-15

3a-9=-15

3a=-15+9

3a=-6

a = -6/3

a = -2

So the solutions are a = -2 and b = 3 giving us the ordered pair (-2,3)